# New Perspectives in Advancing Graph Machine Learning

### v1

@@ Appendix B Proofs of auxiliary lemmas
We first recall the following result from Perlmutter et al. [2023] which shows that the diffusion wavelets $\mathcal{W}_{J}$ are a non-expansive frame on the weighted inner product space defined by $\langle\mathbf{x}_{1},\mathbf{x}_{2}\rangle_{\mathbf{D}^{1/2}}=\langle\mathbf{D}^{1/2}\mathbf{x}_{1},\mathbf{D}^{1/2}\mathbf{x}_{2}\rangle_{2}$, with corresponding norm $\|\mathbf{x}\|_{\mathbf{D}^{1/2}}=\|\mathbf{D}^{1/2}\mathbf{x}\|_{2}$. (See also Proposition 4.1 of Gama et al. [2018] and Theorem 1 of Tong et al. [2024].) The key to the proof of Proposition B.2 will be to combine this result with the inequality (6), stated below, that relates this weighted norm to the unweighted $\ell^{2}$ norm.
Proposition B.1 (Proposition 2.2 of Perlmutter et al. [2023]).
@@
We may use Proposition B.1 to prove the following corollary which analyzes the frame bounds of the diffusion wavelets with respect to the unweighted $\ell^{2}$ norm.
@@
where $d_{\min}$ and $d_{\max}$ denote the minimal and maximal vertex degrees and $c$ is a universal constant.
@@
where $\mathbf{w}[:][k]^{\top}=[\mathbf{w}[1][k],\ldots,x[n][k]]^{\top}$ for all $1\leq k\leq d$. (Formally, let $\boldsymbol{\Pi}$ be the matrix corresponding to the permutation $\widetilde{\sigma}:\{1,2,3,\ldots,nd\}\rightarrow\{1,2,3,\ldots,nd\}$, $\widetilde{\sigma}(i)=n\cdot((i-1)\mod d)+\lceil\frac{j}{d}\rceil$.) To understand this permutation, note that we have reorder the entries of the vector $\mathbf{w}$ so that all entries corresponding to the first output dimension of the function $\mathbf{w}:V\rightarrow\mathbb{R}^{d}$ come first, then all of the entries corresponding to the second output dimension come next, etc.
@@
We first prove the following auxiliary lemma
@@
For $m\geq 1$, we will use induction. In the base case, $m=1$, we will use the definition of block-matrix multiplication, as well as Lemma B.3 to write
$(\overline{\mathbf{Q}}\overline{\mathbf{w}})[i]$ $=(\overline{\mathbf{Q}}(\mathbf{R}\cdot\mathbf{w}))[i]$ (8)
$=\sum_{j=1}^{n}\overline{\mathbf{Q}}[i,j](\mathbf{R}\cdot\mathbf{w})[j]$ (9)
$=\sum_{j=1}^{n}\mathbf{P}[i,j]\mathbf{R}\mathcal{O}_{i,j}\mathbf{R}^{\top}\mathbf{R}\mathbf{w}[j]$ (10)
$=\mathbf{R}\sum_{j=1}^{n}\mathbf{P}[i,j]\mathcal{O}_{i,j}\mathbf{w}[j]$ (11)
$=\mathbf{R}(\mathbf{Q}\mathbf{w}[i]).$ (12)
@@ Appendix C Removing the sign ambiguity
The definition of diffusion wavelets relies on computing the SVD of the matrices $\mathbf{B}_{i}$. However, as noted above, the SVD is only unique up to sign-flips. That is, one may obtain a different SVD $\mathbf{B}^{\prime}_{i}=\mathbf{U}^{\prime}_{i}\boldsymbol{\Sigma}_{i}(\mathbf{V}^{\prime})_{i}^{\top}$ by replacing both $\mathbf{u}_{i,k}$ and $\mathbf{v}_{i,k}$ with $\mathbf{u}^{\prime}_{i,k}=-\mathbf{u}_{i,k}$ and $\mathbf{v}^{\prime}_{i,k}=-\mathbf{v}_{i,k}$ for any fixed $k$. Therefore, we will utilize the sign-flipping trick described below in order to ensure that each of the $\mathcal{O}_{i,j}$ are well defined (and do not suffer from any sign ambiguity).
@@
Therefore, if we are able to ensure that $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is always non-negative, then the definition of $\mathcal{O}_{i,j}$ will not suffer from the sign ambiguity.
Accordingly, when computing $\mathcal{O}_{i,j}$, we check whether or not each $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is positive or negative. If $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is negative, we replace $\mathbf{u}_{j,\ell}$ with $\mathbf{u}_{j,\ell}$. We note that in this construction, the individual vectors, $\mathbf{u}_{i,k}$ still do suffer from the sign-flip ambiguity. However, if one replaces $\mathbf{u}_{i,k}$ with $\mathbf{u}_{i,k}$, the sign flipping trick will also force us to replace $\mathbf{u}_{j,\ell}$ with $\mathbf{u}_{j,\ell}$. Thus, the inner products stay the same since $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle=\langle-\mathbf{u}_{i,k},-\mathbf{u}_{j,\ell}\rangle$.
Appendix D GNN Architecture
In this section, we provide details on the equivariant, scattering-based GNN we used for our experiments in Section 4. (Hyperparameter settings are given in Appendix E.) For simplicity, we will assume that we are given a single vector-valued signal, i.e., $F_{\text{vector}}=1$, which is the case for all of our experiments.
Figure 3: Architecture the equivariant scattering-based GNN. Given a geometric graph, scalar features $\mathbf{X}$ are passed into the scalar track while vector features $\mathbf{W}$ are passed into the vector track. Upon computing the respective scattering coefficients, the updated scalar and vector features are concatenated and projected to form the final node-level embeddings, which can be aggregated to obtain graph-level embeddings.
@@ Appendix D GNN Architecture
We first separately compute the first- and second-order scattering coefficients of each input signal, using (2) for the scalar-valued signals and (4) for the vector-valued signal, and concatenate these to the zeroth-order scattering coefficients (where the zeroth-order coefficients are simply the untransformed signals). We organize the scalar-valued coefficients and vector-valued scattering coefficients into tensors $\widetilde{\mathbf{X}}$ and $\widetilde{\mathbf{W}}$ with dimensions $n\times F_{\text{scalar}}\times S$ and $n\times d\times S$, where $S$ is the total number of zeroth-, first-, and second-order used for each signal at each node. We note that in our experiments, we use the same value of $S$ for all vector- and scalar-valued features. However, we could readily modify our architecture to use a different number of scales for each feature. We illustrate our architecture in Figure 3.
@@
For scalar-valued target functions, we seek for our network to be rationally invariant. Therefore, for node-level tasks, we use a five-layer MLP to map each $\mathbf{t}_{i}$ to the target dimension. Alternatively, for graph-level tasks with scalar targets, we first pool the vectors $\mathbf{t}_{i}$ vectors across nodes within graphs (using, e.g., summation over $i$ or max aggregations over $i$), and again use a five-layer MLP as the final readout head.
@@
We also note that our implementation of the scattering coefficients differs slightly from the exposition in Section 3 for improved performance. Instead of dyadic wavelets of the form $\mathbf{P}^{2^{j-1}}-\mathbf{P}^{2^{j}}$, we instead use generalized diffusion wavelets of the form $\mathbf{P}^{{t_{j-1}}}-\mathbf{P}^{t_{j}}$, where $0=t_{0}<t_{1}<\ldots,<t_{J}$ is an increasing sequence of diffusion scales which are selected by the InfoGain procedure introduced in Johnson et al. [2025b]. Additionally, we did not use a non-linearity while computing the scattering-coefficients since our network is able to learn non-linear relationships in the data via the MLP and the gating function used in (14) and (15). Following Bhaskar et al. [2023], in order to help the geometric scattering transform probe the graph geometry, we also used two Dirac (Kronecker) vectors as additional scalar-valued input signals. To place these Diracs, we first computed the centroid of the vertices in $\mathbb{R}^{d}$ (using the Euclidean distance). We then placed one Dirac at the vertex closest to this centroid and the other Dirac at the vertex furthest from the centroid (ties broken randomly). Importantly, we note that all of our theoretical guarantees may be readily adapted to this modified version of the vector-valued geometric scattering transform.
D.1 Complexity of diffusion wavelets
We compute the (scalar) diffusion matrix $\mathbf{P}$ and the vector diffusion matrix $\mathbf{Q}$ for each graph as a one-time data preprocessing step and cache their (sparse) tensor data in memory. Since operators are independent for each graph, for large data sets, this is an embarrassingly parallel task and so we also batch-parallelize this step across CPU workers.
These constructions of $\mathbf{P}$ and $\mathbf{Q}$ enable highly efficient geometric scattering layers in ESc-GNN, through sparse matrix multiplication of (dense) scalar and vector feature vectors. Still, in terms of complexity of the forward pass of our model, this layer dominates for typical graphs. For each of the scalar and vector tracks, their first-order scattering has complexity $\mathcal{O}\left(S\,m\,(F+d)\right)$, where $S$ is the number of wavelets used (typically four to ten), $F$ is the number of signal channels (generally, $F=1$ for the vector track), and $d$ is the dimension ($d=1$ for scalars). To obtain second-order scattering coefficients, we recursively reapply each (sparse) diffusion operator to all first-order scattering coefficients, and then discard coefficients resulting from higher-pass wavelets reapplied to lower-pass wavelets. Accordingly, for each track, second-order scattering has complexity $\mathcal{O}(n\,F\,d\,S^{2})$, where $n$ is the number of nodes.
Appendix E Experimental details
We use the PyTorch [Ansel et al., 2024] and PyTorch-Geometric [Fey and Lenssen, 2019] to build our codebase, which is available at https://anonymous.4open.science/r/esc-gnn-407D. For the EGNN and TFN models, we adapted code from the “Geometric GNN Dojo” [Joshi et al., 2023] under the MIT license, from their Github repository (https://github.com/chaitjo/geometric-gnn-dojo).
E.1 Cross-validation procedure
@@
E.2 Hyperparameter settings
E.2.1 General
@@
For all models, we optimize for mean squared error (MSE) loss using PyTorch’s AdamW optimizer, an initial learning rate of 0.001, and a train batch size of 32 (except for TFN, for which we use 16, to prevent out-of-memory errors). Models train for at least 50 burn-in epochs, before the following early stopping is enforced: if validation loss does not improve for 50 consecutive epochs (checking every five epochs), we (1) halve the learning rate and reload the best model weights achieved (by by validation loss), and then (2) quit training after a maximum of two such restarts or a maximum of 500 total training epochs. The final model weights are then extracted from the epoch in which the lowest validation loss was achieved.
@@
For experimental consistency and compatibility with our regression tasks, we standardize some modules across models as appropriate. That is, first, for the graph-level diameter estimation task, we use sum and max pooling of hidden node features wherever such pooling is needed; for ESc-GNN, LEGS, GCN, GIN, and GAT, these pooled features are input into a five-layer readout MLP with hidden dimensions 128, 64, 32, and 16. Second, for all models (including ours), we use the sigmoid linear unit (SiLU, or ‘swish’) activation function [Elfwing et al., 2018] in the MLP (except for when another activation function is explicitly prescribed by the baseline model.) Third, when edge weights are used by a model, we compute these weights using a Gaussian kernel $K_{\epsilon}$ from Section 3, with scale parameter $\epsilon$ set equal to the squared mean of mean neighbor distances across the final ellipsoid data set for all models.
E.2.2 ESc-GNN (and ESc-GNN-non-equivariant)
@@
For ESc-GNN, when using the InfoGain Wavelets procedure to select wavelet diffusion scales, we set $t_{J}=16$, and use information cutoff quantiles of $[0.25,0.5,0.75]$ (see Johnson et al. [2025b] for details). We deviated slightly from the method presented in Johnson et al. [2025b] and measured the convergence of each feature as quantified in the $\ell^{1}$ norm rather than the KL divergence. For the scalar signals, we applied the $n\times n$ diffusion matrix $\mathbf{P}$ to each scalar signal $\mathbf{x}$. For the vector signals, we applied the $nd\times nd$ vector diffusion matrix $\mathbf{Q}$ to each vector-valued signal $\mathbf{w}\in\mathbb{R}^{nd}$. This procedure produced diffusion scales $\{t_{j}\}_{j=0}^{J}=\{0,1,2,4,6,8,16\}$ for the scalar features ) and $\{0,1,2,4,6,9,16\}$ for the vector features (after taking the median across input signals as applicable). This means that that our filter bank consists seven total filters (including the low-pass filter $\boldsymbol{\Psi}^{t_{J}}$). In the within-track mixing layers, we set $K$ equal to $16$ and $32$ for the scalar and vector tracks, respectively, and the hidden layer dimensions of their two-layer mixing MLPs to $[64,64]$ and $[128,128]$. (Note the vector track mixing MLP has no biases or activations.) We use sigmoid for the gate activation $\sigma_{w}$.
@@
When ablating the vector track in ESc-GNN to render it non-equivariant, we (1) concatenate the vector feature coordinates to the Dirac scalar features as $d$ separate, additional scalar features, and (2) omit all vector feature operations. (For vector targets, this ablated model then uses the five-layer MLP prediction head defined previously, with an output layer width of the vector dimension $d$.)
E.2.3 LEGS
For LEGS, we use Dirac scalar node features, as in ESc-GNN, and concatenate the vector feature coordinates as $d$ additional scalar features (analogous to ESc-GNN-non-equivariant). The network employs dyadic-scale wavelets with $J=4$. Hidden representations are fed into a five-layer MLP with SiLU activations (and no dropout) to produce predictions. For graph-level predictions, we first pool across nodes using both sum and max operators. For vector targets, the prediction head outputs a vector of dimension $d$.
E.2.4 GCN, GIN, GAT

### v2

@@ Appendix B Proofs of auxiliary lemmas
We first recall the following result from Perlmutter et al. (2023) which shows that the diffusion wavelets $\mathcal{W}_{J}$ are a non-expansive frame on the weighted inner product space defined by $\langle\mathbf{x}_{1},\mathbf{x}_{2}\rangle_{\mathbf{D}^{1/2}}=\langle\mathbf{D}^{1/2}\mathbf{x}_{1},\mathbf{D}^{1/2}\mathbf{x}_{2}\rangle_{2}$, with corresponding norm $\|\mathbf{x}\|_{\mathbf{D}^{1/2}}=\|\mathbf{D}^{1/2}\mathbf{x}\|_{2}$. (See also Proposition 4.1 of Gama et al. (2018) and Theorem 1 of Tong et al. (2024).) The key to the proof of Proposition B.2 is to combine this result with the inequality (9), stated below, that relates this weighted norm to the unweighted $\ell^{2}$ norm.
Proposition B.1 (Proposition 2.2 of Perlmutter et al. (2023)).
@@
We use Proposition B.1 to prove the following corollary which analyzes the frame bounds of the diffusion wavelets with respect to the unweighted $\ell^{2}$ norm.
@@
where $d_{\min}$ and $d_{\max}$ denote the minimal and maximal vertex degrees and $c>0$ is a universal constant.
@@
where $\mathbf{w}[:][k]^{\top}=[\mathbf{w}[1][k],\ldots,\mathbf{w}[n][k]]^{\top}$ for all $1\leq k\leq d$. Formally, let $\boldsymbol{\Pi}$ be the matrix corresponding to the permutation $\widetilde{\sigma}:\{1,2,3,\ldots,nd\}\rightarrow\{1,2,3,\ldots,nd\}$, $\widetilde{\sigma}(i)=n\cdot((i-1)\mod d)+\lceil\frac{j}{d}\rceil$. To understand this permutation, note that we have reorder the entries of the vector $\mathbf{w}$ so that all entries corresponding to the first output dimension of the function $\mathbf{w}:V\rightarrow\mathbb{R}^{d}$ come first, then all of the entries corresponding to the second output dimension come next, etc.
@@
We first prove the following auxiliary lemma.
@@
For $m\geq 1$, we use induction. In the base case, $m=1$, we use the definition of block-matrix multiplication, as well as Lemma B.3 to write
$(\overline{\mathbf{Q}}\overline{\mathbf{w}})[i]$ $=(\overline{\mathbf{Q}}(\mathbf{R}\cdot\mathbf{w}))[i]$
$=\sum_{j=1}^{n}\overline{\mathbf{Q}}[i,j](\mathbf{R}\cdot\mathbf{w})[j]$
$=\sum_{j=1}^{n}\mathbf{P}[i,j]\mathbf{R}\mathcal{O}_{i,j}\mathbf{R}^{\top}\mathbf{R}\mathbf{w}[j]$
$=\mathbf{R}\sum_{j=1}^{n}\mathbf{P}[i,j]\mathcal{O}_{i,j}\mathbf{w}[j]$
$=\mathbf{R}(\mathbf{Q}\mathbf{w}[i]).$
@@ Appendix C Removing the sign ambiguity
The definition of diffusion wavelets relies on computing the SVD of the matrices $\mathbf{B}_{i}$. However, as noted in Section 3, the SVD is only unique up to sign-flips. That is, one may obtain a different SVD $\mathbf{B}^{\prime}_{i}=\mathbf{U}^{\prime}_{i}\boldsymbol{\Sigma}_{i}(\mathbf{V}^{\prime})_{i}^{\top}$ by replacing both $\mathbf{u}_{i,k}$ and $\mathbf{v}_{i,k}$ with $\mathbf{u}^{\prime}_{i,k}=-\mathbf{u}_{i,k}$ and $\mathbf{v}^{\prime}_{i,k}=-\mathbf{v}_{i,k}$ for any fixed $k$. Therefore, we utilize the sign-flipping trick described below in order to ensure that each of the $\mathcal{O}_{i,j}$ are well defined (and do not suffer from any sign ambiguity).
@@
Therefore, if we are able to ensure that $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is always non-negative, then the definition of $\mathcal{O}_{i,j}$ does not suffer from the sign ambiguity.
Accordingly, when computing $\mathcal{O}_{i,j}$, we check whether or not each $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is positive or negative. If $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle$ is negative, we replace $\mathbf{u}_{j,\ell}$ with $\mathbf{u}_{j,\ell}$. We note that in this construction, the individual vectors, $\mathbf{u}_{i,k}$ still do suffer from the sign-flip ambiguity. However, if one replaces $\mathbf{u}_{i,k}$ with $\mathbf{u}_{i,k}$, the sign flipping trick also forces us to replace $\mathbf{u}_{j,\ell}$ with $\mathbf{u}_{j,\ell}$. Thus, the inner products stay the same since $\langle\mathbf{u}_{i,k},\mathbf{u}_{j,\ell}\rangle=\langle-\mathbf{u}_{i,k},-\mathbf{u}_{j,\ell}\rangle$.
Appendix D Complexity of diffusion wavelets
We compute the scalar diffusion matrix $\mathbf{P}$ and the vector diffusion matrix $\mathbf{Q}$ for each graph as a one-time data preprocessing step and cache their (sparse) tensor data in memory. Since operators are independent for each graph, for large data sets this is a parallel task, so we also batch-parallelize this step across CPU workers.
These constructions of $\mathbf{P}$ and $\mathbf{Q}$ enable highly efficient geometric scattering layers in VDW-GNN, through sparse matrix multiplication of (dense) scalar and vector feature vectors. Still, in terms of complexity of the forward pass of our model, this layer dominates for typical graphs. For either a scalar or vector track, first-order scattering has complexity $\mathcal{O}\left(S\,m\,(F+D)\right)$, where $S$ is the number of wavelets used (typically four to ten), $F$ is the number of signal channels (generally, $F=1$ for a vector track), and $D$ is the dimension ($D=1$ for scalars). To obtain second-order scattering coefficients, we recursively reapply each (sparse) diffusion operator to all first-order scattering coefficients, and then discard coefficients resulting from higher-pass wavelets reapplied to lower-pass wavelets. Accordingly, for either track, second-order scattering has complexity $\mathcal{O}(n\,F\,D\,S^{2})$, where $n$ is the number of nodes.
Our codebase and vector diffusion modules are implemented in PyTorch (Ansel et al., 2024), PyTorch-Geometric (Fey and Lenssen, 2019). As such, our method implementations exploit the sparse tensor libraries in these frameworks. Finally, for all experiments, each model training run was done on one NVIDIA L40S 48 GB GPU with CUDA 12.4.
Appendix E Experimental details - Ellipsoids (Section 4.1)
E.1 Architecture
In this section, we provide details on the VDW-GNN model that we used for our ellipsoids experiments in Section 4.1. A graphical illustration of this model is given in Figure 2 of the main text.
@@ E.1 Architecture
We first separately compute the first- and second-order scattering coefficients of each input signal, using (3) and (4) for the scalar-valued signals and (6) and (7) for the vector-valued signal, and concatenate these to the zeroth-order scattering coefficients (where the zeroth-order coefficients are simply the untransformed signals). We organize the scalar-valued coefficients and vector-valued scattering coefficients into tensors $\widetilde{\mathbf{X}}$ and $\widetilde{\mathbf{W}}$ with dimensions $n\times F_{\text{scalar}}\times S$ and $n\times D\times S$, where $S$ is the total number of zeroth-, first-, and second-order used for each signal at each node. (Note that for simplicity, we generally assume that we are given a single vector-valued signal, i.e., $F_{\text{vector}}=1$, which is the case for all of our experiments.) We note that in our experiments, we use the same value of $S$ for all vector- and scalar-valued features. However, we could readily modify our architecture to use a different number of scales for each feature.
@@
For scalar-valued target functions, we design our network to be rotationally invariant. Therefore, for node-level tasks, we use a five-layer MLP to map each $\mathbf{t}_{i}$ to the target dimension. Alternatively, for graph-level tasks with scalar targets, we first pool the vectors $\mathbf{t}_{i}$ vectors across nodes within graphs (using, e.g., summation over $i$ or max aggregations over $i$), and again use a five-layer MLP as the final readout head.
@@
We also note that our implementation of the scattering coefficients differs slightly from the exposition in Section 3 for improved performance. Instead of dyadic wavelets of the form $\mathbf{P}^{2^{j-1}}-\mathbf{P}^{2^{j}}$, we instead use generalized diffusion wavelets of the form $\mathbf{P}^{{t_{j-1}}}-\mathbf{P}^{t_{j}}$, where $0=t_{0}<t_{1}<\ldots,<t_{J}$ is an increasing sequence of diffusion scales which are selected by the InfoGain procedure introduced in Johnson et al. (2025b). Additionally, we did not use a non-linearity while computing the scattering-coefficients since our network is able to learn non-linear relationships in the data via the MLP and the gating function used in (12) and (13). Following Bhaskar et al. (2025), in order to help the geometric scattering transform probe the graph geometry, we also used two Dirac (Kronecker) vectors as additional scalar-valued input signals. To place these Diracs, we first computed the centroid of the vertices in $\mathbb{R}^{d}$ (using the Euclidean distance). We then placed one Dirac at the vertex closest to this centroid and the other Dirac at the vertex furthest from the centroid (ties broken randomly). Importantly, we note that all of our theoretical guarantees may be readily adapted to this modified version of the vector-valued geometric scattering transform.
E.2 Cross-validation procedure
@@
E.3 Hyperparameter settings
E.3.1 General
@@
For all models, we optimize for mean squared error (MSE) loss using PyTorch’s AdamW optimizer, an initial learning rate of 0.001, and a train batch size of 32 (except for TFN, for which we use 16, to prevent out-of-memory errors). Models train for at least 50 burn-in epochs, before the following early stopping is enforced: if validation loss does not improve for 50 consecutive epochs (checking every five epochs), we (1) halve the learning rate and reload the best model weights achieved (by validation loss), and then (2) quit training after a maximum of two such restarts or a maximum of 500 total training epochs. The final model weights are then extracted from the epoch in which the lowest validation loss was achieved.
@@
For experimental consistency and compatibility with our regression tasks, we standardize some modules across models as appropriate. That is, first, for the graph-level diameter estimation task, we use sum and max pooling of hidden node features wherever such pooling is needed; for VDW-GNN, LEGS, GCN, GIN, and GAT, these pooled features are input into a five-layer readout MLP with hidden dimensions 128, 64, 32, and 16. Second, for all models (including ours), we use the sigmoid linear unit (SiLU, or ‘swish’) activation function (Elfwing et al., 2018) in the MLP (except for when another activation function is explicitly prescribed by the baseline model.) Third, when edge weights are used by a model, we compute these weights using a Gaussian kernel $K_{\epsilon}$ from Section 3, with scale parameter $\epsilon$ set equal to the squared mean of mean neighbor distances across the final ellipsoid data set for all models.
E.3.2 VDW-GNN (and VDW-GNN-non-equivariant)
@@
For VDW-GNN, when using the InfoGain Wavelets procedure to select wavelet diffusion scales, we set $t_{J}=16$, and use information cutoff quantiles of $[0.25,0.5,0.75]$ (see Johnson et al. (2025b) for details). We deviated slightly from the method presented in Johnson et al. (2025b) and measured the convergence of each feature as quantified in the $\ell^{1}$ norm rather than the KL divergence. For the scalar signals, we applied the $n\times n$ diffusion matrix $\mathbf{P}$ to each scalar signal $\mathbf{x}$. For the vector signals, we applied the $nD\times nD$ vector diffusion matrix $\mathbf{Q}$ to each vector-valued signal $\mathbf{w}\in\mathbb{R}^{nd}$. This procedure produced diffusion scales $\{t_{j}\}_{j=0}^{J}=\{0,1,2,4,6,8,16\}$ for the scalar features ) and $\{0,1,2,4,6,9,16\}$ for the vector features (after taking the median across input signals as applicable). This means that that our filter bank consists seven total filters (including the low-pass filter $\boldsymbol{\Psi}^{t_{J}}$). In the within-track mixing layers, we set $K$ equal to $16$ and $32$ for the scalar and vector tracks, respectively, and the hidden layer dimensions of their two-layer mixing MLPs to $[64,64]$ and $[128,128]$. (Note the vector track mixing MLP has no biases or activations.) We use sigmoid for the gate activation $\sigma_{w}$.
@@
When ablating the vector track in VDW-GNN to render it non-equivariant, we (1) concatenate the vector feature coordinates to the Dirac scalar features as $D$ separate, additional scalar features, and (2) omit all vector feature operations. (For vector targets, this ablated model then uses the five-layer MLP prediction head defined previously, with an output layer width of the vector dimension $D$.)
E.3.3 LEGS
For LEGS, we use Dirac scalar node features, as in VDW-GNN, and concatenate the vector feature coordinates as $d$ additional scalar features (analogous to VDW-GNN-non-equivariant). The network employs dyadic-scale wavelets with $J=4$. Hidden representations are fed into a five-layer MLP with SiLU activations (and no dropout) to produce predictions. For graph-level predictions, we first pool across nodes using both sum and max operators. For vector targets, the prediction head outputs a vector of dimension $d$.
E.3.4 GCN, GIN, GAT